OpenStudy (lgbasallote):

Openstudy user L is level 70 and increases levels once in two days. Openstudy user P is level 1 and increases one and a half level every day. How many days will it take for OS user P to overtake OS user L?

Parth (parthkohli):

It's you again -.-

OpenStudy (lgbasallote):

hi again

Parth (parthkohli):

lol P = Parth and L=Lgba you cheat xP

OpenStudy (lgbasallote):

lol you catch on fast

OpenStudy (lgbasallote):

it came from my book though

Parth (parthkohli):

I'm better than you.. get it? B)

OpenStudy (anonymous):

xD you guys are awesome. I love the competitive spirit xD

Parth (parthkohli):

Lgba, you were 86 when I joined.

Parth (parthkohli):

Yet, I caught you once in the race xP

Parth (parthkohli):

Remember the horror on your face when I was ahead? lol

Parth (parthkohli):

They'll catch up on 99(both)

OpenStudy (lgbasallote):

lol you caught up when we were 99 :P

OpenStudy (lgbasallote):

and this is a serious question

Parth (parthkohli):

Yeah.. they can't catch up earlier than that

OpenStudy (lgbasallote):

well assuming there are infinite levels?

Parth (parthkohli):

3 levels every two days for P. 1 level every two days for L.

Parth (parthkohli):

Let's assume they'll meet at \(x\). what will be the equation?

Parth (parthkohli):

Not sure what to do now :P

Parth (parthkohli):

@lgbasallote did you understand? can you solve now?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x = 70 + 0.5x }\) Something like that?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2.5x = 70 }\) \( \color{Black}{\Rightarrow x = 28 }\) 28 days!

Parth (parthkohli):

lol but it took me 40 days xP

Parth (parthkohli):

@lgbasallote you there?

OpenStudy (lgbasallote):

in 28 days P is lvl 42

Parth (parthkohli):

So like 112 levels.

OpenStudy (lgbasallote):

they dont meet in 28 days

Parth (parthkohli):

Oh :/

Parth (parthkohli):

then I suck at math lol

OpenStudy (lgbasallote):

@mathslover help?

OpenStudy (lgbasallote):

im half-afk so bear with me if i dont reply for a long time

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x = 70 + 0.5x }\) I'm trying to find the mistake :/

mathslover (mathslover):

wait

OpenStudy (turingtest):

1+2x=70+0.5x ????

OpenStudy (lgbasallote):

explanation for equation?

Parth (parthkohli):

One and a half. umm..

OpenStudy (lgbasallote):

this is an even number in my book so no answer :(

OpenStudy (turingtest):

oh, then 1+1.5x=70+0.5x

Parth (parthkohli):

Oh yes.. forgot that 1

Parth (parthkohli):

\( \color{Black}{\Rightarrow1.5x = 69 + 0.5x }\) \( \color{Black}{\Rightarrow x = 69 }\) lol why did I put that 3x lol

OpenStudy (lgbasallote):

what's that 1 for?

Parth (parthkohli):

I'm verified with the idiot stamp

Parth (parthkohli):

I start at level 1,

Parth (parthkohli):

Though, I reached 57 on the first day =)

OpenStudy (lgbasallote):

lol this aint you boy...so explanation for equation?

OpenStudy (turingtest):

yeah, I have a few accounts the smartscore can go up much faster than the old system

OpenStudy (turingtest):

parth do you want to explain it?

Parth (parthkohli):

Yes.

Parth (parthkohli):

Because I start at level 1 lolol

Parth (parthkohli):

And 1.5x means my increase,

Parth (parthkohli):

As lgba increases one in two days, it can be said that it's half per day

OpenStudy (lgbasallote):

why +?

Parth (parthkohli):

Lgba starts with 70, so that's the reason we have the 70

OpenStudy (lgbasallote):

why +?

Parth (parthkohli):

70 + your rate of increase, because your levels won't decrease lol

OpenStudy (lgbasallote):

what is this formula o.O

Parth (parthkohli):

70 is the initial value.

OpenStudy (lgbasallote):

what's x? maybe it can make sense

OpenStudy (turingtest):

\[\text{current level}=\text{starting level}+\text{rate level change}\times\text{time}\]

Parth (parthkohli):

You're adding the rate of change in the initial value. Similarly, my initial value is 1.

OpenStudy (lgbasallote):

oh i see

OpenStudy (lgbasallote):

now itamakes sens

Parth (parthkohli):

x is the number of days

OpenStudy (lgbasallote):

darnkeyboard

Parth (parthkohli):

lol

OpenStudy (turingtest):

that's what I meant :P^

Parth (parthkohli):

We have the rate of change, so the variable would be the time.

OpenStudy (lgbasallote):

i get it now thanks

Parth (parthkohli):

yw, son

Parth (parthkohli):

;p

Parth (parthkohli):

A guy who knows calculus, but is confused in the algebra problem..

OpenStudy (lgbasallote):

we all have heels...the world is round after all

Parth (parthkohli):

Well of course Idk cal but I know algebra lol

OpenStudy (turingtest):

semi-dyslexia and absent-mindedness persist no matter what you learn ;)

Parth (parthkohli):

^ lgba that goes for you;p

OpenStudy (lgbasallote):

my semi-dyslexia refuses to understand what that meant

OpenStudy (lgbasallote):

there was a follow up question....suppose P stopped 10 days after he joined OS and remained at rest for 3 days how long will it take for them to equal now?

Parth (parthkohli):

omg

OpenStudy (lgbasallote):

blame the book

OpenStudy (lgbasallote):

imnot making these questions

Parth (parthkohli):

Y U NO KNOW ALGEBRA

OpenStudy (lgbasallote):

coz i hate physics...and this is related to that

OpenStudy (lgbasallote):

how do i set upthe equation?

Parth (parthkohli):

Where's physics here

OpenStudy (lgbasallote):

speed time and distane lol

Parth (parthkohli):

First figure out what the level of lgba would be after 13 days, that'd be easy :D

Parth (parthkohli):

76.5

Parth (parthkohli):

Now set up the same equation

OpenStudy (lgbasallote):

i see thanks

Parth (parthkohli):

yw, son

OpenStudy (lgbasallote):

70.5 days?

OpenStudy (anonymous):

@.@